OpenStudy (anonymous):

determine the slope and y-intercept for the equation 6x+3y=9

OpenStudy (anonymous):

First, isolate y.

OpenStudy (anonymous):

3y = 9 - 6x Divide everything by 3 y = 3 - 2x Replace y with f(x) f(x) = 3 - 2x

OpenStudy (anonymous):

f(x) = -2x + 3

OpenStudy (anonymous):

To find the y intercept, let x = 0.

OpenStudy (anonymous):

f(0) = -2(0) + 3 f(0) = 3 Remember that y = f(x) y = 3

OpenStudy (anonymous):

So when x is 0, y is 3. therefor, (0,3) is your y intercept.

OpenStudy (anonymous):

To calculate the slope, we need another ordered pair. Let x = 5 f(5) = -2(5) + 3 f(5) = -10 + 3 f(5) = -7 y = -7 x = 5 y = -7 (5,-7)

OpenStudy (anonymous):

Next, we need to use the slope formula.

OpenStudy (anonymous):

y2 - y1 over x2 - x1

OpenStudy (anonymous):

-7 - 3 over 5 - 0 -10 over 5 -2 So your slope is -2, and your y intercept is (0,3).

OpenStudy (anonymous):

Welcome to OpenStudy. :D

OpenStudy (goformit100):

A Warm Welcome To Open Study